OpenStudy (anonymous):

i was using during the finding area of circle that is 22/7*28 square=22/7*784 then we get different answer than if we use 3.14*784.what's the logic behind it.can anyone tell me ...

OpenStudy (cggurumanjunath):

radius =28 right ?

OpenStudy (cggurumanjunath):

@ziqbal103

OpenStudy (anonymous):

yes

OpenStudy (cggurumanjunath):

2464(area) when 22/7 =pi 2461.76(area) when pi =3.14 right ?

OpenStudy (cggurumanjunath):

22/7=3.142857

OpenStudy (cggurumanjunath):

pi=3.14

OpenStudy (cggurumanjunath):

https://www.google.co.in/#q=22%2F7

OpenStudy (cggurumanjunath):

answer is different because value of pi inputted is different.

OpenStudy (cggurumanjunath):

@ziqbal103 is it clear now ?

OpenStudy (anonymous):

yes when we use 22/7 then we get answer 2646, but when we use 3.14 *28square then we don't get the same answer

OpenStudy (anonymous):

but what answer should we prefer in some questions of oxford book they are using the value of pi 22/7 and in some questions 3.14.

OpenStudy (cggurumanjunath):

very good.

OpenStudy (cggurumanjunath):

if the value of pi is given as take pi=3.14 then use pi=3.14 in your answer if the value of pi is given as take pi=22/7 then use pi=22/7 in your answer

OpenStudy (cggurumanjunath):

it it it said round your answer to 2 decimal digits then take pi=3.14

OpenStudy (cggurumanjunath):

http://hotmath.com/help/gt/genericprealg/section_9_3.html

OpenStudy (cggurumanjunath):

go through the above useful link to understand better.thank you.

OpenStudy (anonymous):

thank you very muuch

OpenStudy (cggurumanjunath):

wc.